OpenStudy (anonymous):

I have a problem which is asking for amplitude and the period and to sketch th graph. How would i do y = 4 sin x Please show some work so i can learn how to do this to finish the rest.

OpenStudy (amistre64):

amplitude is the scalar; akin to slope

OpenStudy (amistre64):

period is normal in this case; as is all the rest in it

OpenStudy (phi):

Here's a plot http://www.wolframalpha.com/input/?i=plot++y+%3D+4+sin+x

OpenStudy (phi):

The amplitude is the height of the sin wave above zero. It's the 4 in front of the sin The period is one wave length. one way to find it is remember (memorize) in \(\sin( \frac{2\pi}{T} x)\) T is the period. so in your case set up an equation between the coefficient of x in sin(x) = 2pi/T Here the coefficient is just 1 so 1= 2pi/T and T= 2 pi

OpenStudy (phi):

If you had sin(2x) the amplitude is 1 to find the period: solve 2 = 2 pi/T for T: T= pi